# Reevaluating Policy Gradient Methods for Imperfect-Information Games

### v1

@@ A.2 State-of-the-Art Tabular Solvers
    CFR+ (cfr+14) subtitutes the regret matching (RM) algorithm in CFR with regret matching+ (RM+), leading to better performance empirically. The main difference between RM and RM+ is that RM+ clamps the regrets to be non-negative.
@@
    Discounted CFR (DCFR) (brown2019solving), in a similar vain as CFR+, discounts the regrets to reduce the effect of previous actions with large (absolute) regret. Practically speaking, at iteration $t$, we rescale the negative regrets by $1-1/(1+t^{\beta})$ and the positive regrets by $1-1/(1+t^{\alpha})$. We use the default value of $\alpha=1.5$ and $\beta=0$.
@@
    Predictive CFR (PCFR) (farina2019stable) uses optimistic regret minimizers to achieve empirical and theoretical faster convergence rates for CFR. Practically, we observe the gradients twice to calculate the behavior policy of CFR.
@@
[…]
@@ Appendix B Solving Dark Hex 3
We investigate a deterministic winning strategy for the first moving player in Dark Hex 3. The value of this policy against a uniform policy is 1, implying that it wins against all possible deterministic strategies, including any possible best response to it. The strategy is shown in Figure 6 where, at each information state, a list of actions to try in that order is given. If the first action fails, the next action is played, and so on. After playing an action, there are two possible outcomes: the action is playable, or information is gained. In general, the size of the list of actions is equal to one plus the number of opponent pieces that have been played but have not been observed; thus, the last action has to be playable. Beyond the computational proof that the strategy in Figure 6 is optimal, we can show it using the fact the game of Hex (with perfect information) cannot end in a draw (gale1979game, Hex Theorem). Since the first player always wins in the reachable information states, it is impossible the invisible pieces are in a way that the opponent has already won.
@@ Appendix C Algorithm implementation details
(Source Implementation): We use the implementation of NFSP in OpenSpiel (openspiel19). This implementation learns distinct models for player 1 and player 2. For consistency with the other algorithms, we rewrote some of the algorithm to use PyTorch (pytorch19) instead of TensorFlow (tf15).
@@
(Source Implementation): We use the implementation of PSRO in OpenSpiel (openspiel19). This implementation learns distinct models for player 1 and player 2. For consistency with the other algorithms, we replaced the default OpenSpiel Tensorflow DQN agent (source) with the OpenSpiel PyTorch DQN agent (source).
@@
(Source Implementation): We use the implementation of R-NaD in OpenSpiel (openspiel19). This implementation learns a single model that is used for both player 1 and player 2. Natively, this algorithm is written in JAX (jax2018github). For convenience during evaluation, we convert the neural networks learned in JAX to equivalent PyTorch models. We verified this conversion by ensuring that, for the same inputs, the outputs of the models are equivalent.
@@
(Source Implementation): We use the implementation of ESCHER from the original paper (escher23). This implementation learns a single policy network that is used for both player 1 and player 2. Natively this algorithm is written in Tensorflow. For convenience during evaluation we convert the neural networks learned in Tensorflow to equivalent PyTorch models and verify by ensuring each layer weight and bias is equivalent.
@@ D.2 MMD
We verified that we obtain results consistent with the original MMD implementation (mmd23). Namely, after a 10M steps training we obtain an approximate exploitability of around 0.14 and 0.20 on PTTT and ADH3 respectively. For reference, Table 2 in mmd23 reports an approximate exploitability of $0.15\pm 0.01$ and $0.20\pm 0.01$ for PTTT and ADH3 respectively.
Appendix E Additional results
@@
Figure 10: The top 10 most influencial hyperparameters from the hyperparameter tuning launch, broken down by game and algorithm. The sweep results provide mappings from hyperparameter sets to exploitabilities, which we use to train a random forest regression model. This model assigns each hyperparameter a coefficient representing its importance in the prediction. This importance reflects the hyperparameter’s impact on exploitability: a high value indicates a strong influence, while a low value suggests minimal impact.
@@
Library Entropy Coefficient
Stable Baselines (hill2018stablebaselines) 0.0
CleanRL (cleanrl22) 0.01 or 0.001
RLlib (liang2018rllib) 0.0
OpenSpiel (openspiel19) 0.01
PufferLib (suarez2024pufferlibmakingreinforcementlearning) 0.01
RL-Games (rl-games2021) 0 or 0.01
Tianshou (tianshou) 0 or 0.01
@@
Appendix F Algorithm hyperparameters
@@ F.2 PSRO
    • Description: Learning rate for Q-network.
    • Default value: 0.010.01 0.010.01
@@ F.3 R-NaD
    • Description: Number of samples in each training batch.
    • Default value: 256256 256256
@@
    • Description: Coefficient for V-trace importance weights (impala18).
    • Default value: 1.01.0 1.01.0
@@
    • Description: Size of the gradient clipped threshold in the NeurD gradient clipping (hennes2019neural).
    • Default value: 2.02.0 2.02.0
@@
    • Description: Size of clipping for the importance sampling in the NeurD gradient clipping (hennes2019neural).
@@ F.4 ESCHER
    • Source: Hyperparameters for Phantom TTT and Dark Hex 4 from escher23.
@@
    • Source: Hyperparameters for Phantom TTT and Dark Hex 4 from escher23.
@@
    • Source: Hyperparameters for Phantom TTT and Dark Hex 4 from escher23.
@@
    • Source: Hyperparameters for Phantom TTT and Dark Hex 4 from escher23.
@@
    • Source: Hyperparameters for Phantom TTT and Dark Hex 4 from escher23.
@@
    • Source: Hyperparameters for Phantom TTT and Dark Hex 4 from escher23.
@@
    • Source: Hyperparameters for Phantom TTT and Dark Hex 4 from escher23.
@@ F.5 PPO
    • Description: The number of minibatches (i.e. the minibatch size is round(num_steps x num_envs / num_minibatches).
    • Default value: 44 44
@@
    • Description: Number of policy update epochs (i.e. how many times to go through the whole batch in each iteration).
    • Default value: 44 44
@@
    • Default value: 0.050.05 0.050.05
    • Source: Inspired from (mmd23).
    • Note: This value is larger than in OpenSpiel’s PPO Implementation because we have found that entropy bonuses lead to much better policies. However, the default value of 0.010.01 0.010.01 is still within our sweeping range.
@@
F.6 PPG
We consider the same parameters as in PPO (Section F.5), with the addition of the following ones:
@@ F.6 PPG
    • Default value: 3232 3232
    • Source: CleanRL implementation and ppg20.
@@
    • Default value: 11 11
    • Source: CleanRL implementation and ppg20.
@@
    • Default value: 11 11
    • Source: CleanRL implementation and ppg20.
@@
    • Default value: 66 66
    • Source: CleanRL implementation and ppg20.
@@
    • Default value: 1.01.0 1.01.0
    • Source: CleanRL implementation and ppg20.
@@
    • Default value: 11 11
F.7 MMD
We consider the same parameters as in PPO (Section F.5), with the addition of the following ones:
@@ F.7 MMD
    • Default value: 0.050.05 0.050.05
    • Source: (mmd23).

### v2

@@ A.2 State-of-the-Art Tabular Solvers
    CFR+ [Tammelin, 2014] substitutes the regret matching (RM) algorithm in CFR with regret matching+ (RM+), leading to better performance empirically. The main difference between RM and RM+ is that RM+ clamps the regrets to be non-negative.
@@
    Discounted CFR (DCFR) [Brown and Sandholm, 2019b], in a similar vain as CFR+, discounts the regrets to reduce the effect of previous actions with large (absolute) regret. Practically speaking, at iteration $t$, we rescale the negative regrets by $1-1/(1+t^{\beta})$ and the positive regrets by $1-1/(1+t^{\alpha})$. We use the default value of $\alpha=1.5$ and $\beta=0$.
@@
    Predictive CFR (PCFR) [Farina et al., 2019] uses optimistic regret minimizers to achieve empirical and theoretical faster convergence rates for CFR. Practically, we observe the gradients twice to calculate the behavior policy of CFR.
@@
Game Algorithm Value Exploitability
CFR 0.9999993 0.0000004
CFR+ 0.9999993 0.0000004
DCFR 0.9999993 0.0000004
FP 0.9999997 0.0000002
PCFR 0.9999993 0.0000005
PCFR+ 0.9999993 0.0000005
3x3 Dark Hex (DH3) PDCFR 0.9999993 0.0000005
CFR 0.3841795 0.0295619
CFR+ 0.3844054 0.0011629
DCFR 0.3844321 0.0010063
FP 0.3580483 0.1929647
PCFR 0.3844107 0.0011789
PCFR+ 0.3843936 0.0006123
3x3 Abrupt Dark Hex (ADH3) PDCFR 0.3844073 0.0192829
CFR 0.6664844 0.0107796
CFR+ 0.6666324 0.0016106
DCFR 0.6666511 0.0004070
FP 0.6622969 0.1016475
PCFR 0.6660216 0.0011638
PCFR+ 0.6662418 0.0019337
Phantom Tic-Tac-Toe (PTTT) PDCFR 0.6662910 0.0157007
CFR 0.5501850 0.0162306
CFR+ 0.5501825 0.0004936
DCFR 0.5501728 0.0003577
FP 0.5428041 0.2871734
PCFR 0.5501960 0.0011708
PCFR+ 0.5501886 0.0004846
Abrupt Phantom Tic-Tac-Toe (APTTT) PDCFR 0.5502121 0.0040427
@@ Appendix B Solving Dark Hex 3
We investigate a deterministic winning strategy for the first moving player in Dark Hex 3. The value of this policy against a uniform policy is 1, implying that it wins against all possible deterministic strategies, including any possible best response to it. The strategy is shown in Figure 5 where, at each information state, a list of actions to try in that order is given. If the first action fails, the next action is played, and so on. After playing an action, there are two possible outcomes: the action is playable, or information is gained. In general, the size of the list of actions is equal to one plus the number of opponent pieces that have been played but have not been observed; thus, the last action has to be playable. Beyond the computational proof that the strategy in Figure 5 is optimal, we can show it using the fact the game of Hex (with perfect information) cannot end in a draw [Gale, 1979, Hex Theorem]. Since the first player always wins in the reachable information states, it is impossible the invisible pieces are in a way that the opponent has already won.
@@ Appendix C Algorithm implementation details
(Source Implementation): We use the implementation of NFSP in OpenSpiel [Lanctot et al., 2019]. This implementation learns distinct models for player 1 and player 2. For consistency with the other algorithms, we rewrote some of the algorithm to use PyTorch [Paszke et al., 2019] instead of TensorFlow [Abadi et al., 2015].
@@
(Source Implementation): We use the implementation of PSRO in OpenSpiel [Lanctot et al., 2019]. This implementation learns distinct models for player 1 and player 2. For consistency with the other algorithms, we replaced the default OpenSpiel Tensorflow DQN agent (source) with the OpenSpiel PyTorch DQN agent (source).
@@
(Source Implementation): We use the implementation of R-NaD in OpenSpiel [Lanctot et al., 2019]. This implementation learns a single model that is used for both player 1 and player 2. Natively, this algorithm is written in JAX [Bradbury et al., 2018]. For convenience during evaluation, we convert the neural networks learned in JAX to equivalent PyTorch models. We verified this conversion by ensuring that, for the same inputs, the outputs of the models are equivalent.
@@
(Source Implementation): We use the implementation of ESCHER from the original paper [McAleer et al., 2023]. This implementation learns a single policy network that is used for both player 1 and player 2. Natively this algorithm is written in Tensorflow. For convenience during evaluation we convert the neural networks learned in Tensorflow to equivalent PyTorch models and verify by ensuring each layer weight and bias is equivalent.
@@ D.2 MMD
We verified that we obtain results consistent with the original MMD implementation [Sokota et al., 2023]. Namely, after a 10M steps training we obtain an approximate exploitability of around 0.14 and 0.20 on PTTT and ADH3 respectively. For reference, Table 2 in Sokota et al. [2023] reports an approximate exploitability of $0.15\pm 0.01$ and $0.20\pm 0.01$ for PTTT and ADH3 respectively.
Appendix E Entropy Coefficient Analysis
Library Entropy Coefficient
Stable Baselines [Hill et al., 2018] 0.0
CleanRL [Huang et al., 2022] 0.01 or 0.001
RLlib [Liang et al., 2018] 0.0
OpenSpiel [Lanctot et al., 2019] 0.01
PufferLib [Suarez, 2024] 0.01
RL-Games [Makoviichuk and Makoviychuk, 2021] 0 or 0.01
Tianshou [Weng et al., 2022] 0 or 0.01
In Section 3, we suggest that one possible reason for the poor performance of generic PG methods in existing literature is that the entropy coefficients required to achieve good performance are too far removed from typical choices. To investigate, we plot exploitability as a function of entropy coefficient in Figure 7. The results are consistent with this possible explanation: Not only do unilateral changes in entropy coefficient produce drastic changes in exploitability, but the entropy coefficients with the best average performance are between 0.05 and 0.2, larger than any of the default entropy coefficients for PPO in Stable Baselines [Hill et al., 2018], CleanRL [Huang et al., 2022], RLlib [Liang et al., 2018], OpenSpiel [Lanctot et al., 2019], PufferLib [Suarez, 2024], RL-Games [Makoviichuk and Makoviychuk, 2021], and Tianshou [Weng et al., 2022], which, as detailed in Table 3, range between 0 and 0.01.
Appendix F Additional results
@@
Figure 10: The top 10 most influencial hyperparameters from the hyperparameter tuning launch, broken down by game and algorithm. The sweep results provide mappings from hyperparameter sets to exploitabilities, which we use to train a random forest regression model. This model assigns each hyperparameter a coefficient representing its importance in the prediction. This importance reflects the hyperparameter’s impact on exploitability: a high value indicates a strong influence, while a low value suggests minimal impact. Note: inner_rl_agent was abbreviated to inner.
@@
Appendix G Algorithm hyperparameters
@@ G.3 R-NaD
    • Description: Size of the gradient clipped threshold in the NeurD gradient clipping [Hennes et al., 2019].
    • Default value: $2.0$
@@
    • Description: Size of clipping for the importance sampling in the NeurD gradient clipping [Hennes et al., 2019].
@@ G.4 ESCHER
    • Source: Hyperparameters for Phantom TTT and Dark Hex 4 from McAleer et al. [2023].
@@
    • Source: Hyperparameters for Phantom TTT and Dark Hex 4 from McAleer et al. [2023].
@@
    • Source: Hyperparameters for Phantom TTT and Dark Hex 4 from McAleer et al. [2023].
@@
    • Source: Hyperparameters for Phantom TTT and Dark Hex 4 from McAleer et al. [2023].
@@
    • Source: Hyperparameters for Phantom TTT and Dark Hex 4 from McAleer et al. [2023].
@@
    • Source: Hyperparameters for Phantom TTT and Dark Hex 4 from McAleer et al. [2023].
@@
    • Source: Hyperparameters for Phantom TTT and Dark Hex 4 from McAleer et al. [2023].
@@ G.5 PPO
    • Default value: $0.05$
    • Source: Inspired from [Sokota et al., 2023].
    • Note: This value is larger than in OpenSpiel’s PPO Implementation because we have found that entropy bonuses lead to much better policies. However, the default value of $0.01$ is still within our sweeping range.
@@
G.6 PPG
We consider the same parameters as in PPO (Section G.5), with the addition of the following ones:
@@ G.6 PPG
    • Default value: $32$
    • Source: CleanRL implementation and Cobbe et al. [2021].
@@
    • Default value: $1$
    • Source: CleanRL implementation and Cobbe et al. [2021].
@@
    • Default value: $1$
    • Source: CleanRL implementation and Cobbe et al. [2021].
@@
    • Default value: $6$
    • Source: CleanRL implementation and Cobbe et al. [2021].
@@
    • Default value: $1.0$
    • Source: CleanRL implementation and Cobbe et al. [2021].
@@
    • Default value: $1$
G.7 MMD
We consider the same parameters as in PPO (Section G.5), with the addition of the following ones:
@@ G.7 MMD
    • Default value: $0.05$
    • Source: [Sokota et al., 2023].